OpenStudy (anonymous):

with non uniform circular motion to find the acceleration of R do you do V^2/magnitude of R

OpenStudy (turingtest):

what do you mean the acceleration of R ? could you elaborate?

OpenStudy (anonymous):

well i want to find the tension of the rope of the pendulum and to do that i have to take ma and im given the mass of the pendulum and length, but not the acceleration

OpenStudy (anonymous):

and im given the velocity

OpenStudy (turingtest):

does it say at what point it is in the swing?

OpenStudy (anonymous):

ya as its swinging upward and it makes a certain angle with the vertical

OpenStudy (turingtest):

all the forces along the r direction must be zero, so let's consider that

OpenStudy (anonymous):

yep looks like what i drew

OpenStudy (turingtest):

so along r we have the force of gravity in that direction and the centripetal force being cancelled by tension. Notice that the component of gravity along r will be the cosine of the angle that the string makes with the vertical:\[\large \sum F_r=F_{gr}+F_c-T=mg\cos\theta+m\frac{v^2}r-T=0\]this can be solved for the tension T